OpenStudy (anonymous):

a 100lb block of uranium is decaying into lead exponentially using the equation Q(t)=Q e^kt. if 85 lbs remains after 5 years how many years will there be 10 lbs or uranium?

OpenStudy (nottim):

is this a plug and play situation?

OpenStudy (nottim):

i havent really read the question, but that's my first guess.

jimthompson5910 (jim_thompson5910):

Q(t)=Q*e^(kt) Q(t)=100*e^(kt) 85=100*e^(k*5) 85/100=e^(k*5) 0.85=e^(k*5) ln(0.85)=5k ln(0.85)/5=k -0.0325 = k k = -0.0325 ------------------------------------------------------- Q(t)=100*e^(kt) Q(t)=100*e^(-0.0325t) 10=100*e^(-0.0325t) 10/100=e^(-0.0325t) 0.1=e^(-0.0325t) ln(0.1)=-0.0325t ln(0.1)/(-0.0325) = t 70.8488= t t = 70.8488 So in approximately 70.8488 years, there will be 10 lbs of uranium left.